Please help me write out and solve this equation:
A store owener sells two pairs of running shoes for each pair of high tops. He plans to order 144 pairs of shoes. How many pairs should be high tops?

Found 2 solutions by unlockmath, scott8148:
Answer by unlockmath(1122)

Hello,
Let's first set up where x will show the number of high tops. 2x will represent the number of running shoes. Now we can set up and equation:
x+2x=144 Combine like terms
3x=144 Divide both sides by 3
x=48
This shoes us that 48 high tops will be sold and 96 Running shoes will be sold.
RJ Toftness
www.math-unlock.com

Answer by scott8148(5892)

"A store owener sells two pairs of running shoes for each pair of high tops" ___ r = 2h

"He plans to order 144 pairs of shoes" ___ r + h = 144

two unknowns (r and h) and two equations

substituting ___ (2h) + h = 144 ___ h = 48

substituting ___ r = 2(48) = 96